OpenStudy (anonymous):

Solve the system 9=4x-2y 2x+6=y

OpenStudy (johnweldon1993):

Well for this one...you already know what y equals via that second equation y = 2x + 6 so you can substitute that in for 'y' in the first equation 9 = 4x - 2y 9 = 4x - 2(2x + 6) Now can you solve for 'x' ...?

OpenStudy (anonymous):

no how do you solve for x?

OpenStudy (johnweldon1993):

We would simplify that equation 9 = 4x - 2(2x + 6) 9 = 4x - 4x - 12 9 = -12 there is no answer to this system of equations Can anyone make sure I did that right? not too often is there no solution...

OpenStudy (anonymous):

matrices says your right john!

OpenStudy (johnweldon1993):

I really should learn how to solve these via matrices...never learned that method Thanks @chris00

OpenStudy (anonymous):

well you can simply see that their is no solution by inspection!

OpenStudy (anonymous):

the second equation is a scalar multiple of the first equation, however their constant's are not.

OpenStudy (anonymous):

so since they are scalar multiples, then their can either be parallel and distinct ( don't lie on the same line) or they are parallel and coincident (they are vertually the same line...

OpenStudy (anonymous):

since 6 Is not a scalar multiple of 9 in this case then the system is inconsistent based on the fact that these lines are parallel and distinct.